# Inference Optimal VLMs Need Only One Visual Token but Larger Models

### v1

@@ 3.3.2 Variation in Optimal Tokens with Text Query Length
(a) Performance trends and trade-offs of VLMs change when varying the number of input text token $Q$.
(b) Scaling laws on OCR-like tasks favor visual token count over LLM size; the opposite of visual reasoning.
@@
Figure 3: Adjusting input text token count and benchmark family shifts performance trends. Left (a): For visual reasoning tasks, as the number of text tokens $Q$ increases, the impact of increasing the number of visual tokens $V$, i.e., reducing compression, becomes more apparent. Intuitively, at enough text tokens, initial increases in visual tokens are only a minor fraction of the overall compute (§ 3.3.2). Right (b): When tasks are changed from visual reasoning to OCR/text-understanding, trends reverse: visual token count should now be prioritized over LLM size (§ 3.3.3).
The shift in performance trends and the ideal visual token count from $Q=0\rightarrow 50$ raises the question; how does the input text length impact the optimal selection of LLM size and number of visual tokens? To explore the variations in trends, we consider the effect of text input length on the optimal inference behavior in Figure 3(a). First, when the text input length, $Q$, is small (purple curves), it is always better to use the larger model (solid curve) with less visual tokens compared to the smaller model (dashed curve) with more visual tokens. However, consider an edge case where the text input length is extremely high (e.g., 100 for the green curves). We observe that there is a sharp increase in error as inference FLOPs are reduced. This is because visual tokens need to be reduced significantly for any effective change in inference FLOPs, as the fixed cost from text tokens is quite high. At a certain point (marked by the red dot in Figure 3(a)), it becomes more advantageous to use the 4B model with a higher number of visual tokens rather than the 7B model with fewer tokens (contrary to the case for lower $Q$). Thus, the optimal number of visual input tokens rises with an increase in $Q$. This case demonstrates the need for careful balancing of visual token count and LLM size, especially in scenarios where text inputs are long, to achieve compute-optimal performance without sacrificing accuracy.
Despite the changes in the exact optimal visual token count and LLM parameter count as the length of the user query increases, the general trend for visual reasoning and understanding tasks is that increasing the size of the language model while reducing visual tokens can lead to significant relative gains (as also illustrated in Fig. 4). This finding may be due, in part, to the scaling properties of the LLMs, which allow larger models to extrapolate with less visual information than their smaller counterparts (Radford et al., 2021; Wei et al., 2022). However, this trade-off does not extend to certain tasks, such as document comprehension, text identification, etc., where a single or handful of tokens may not be able to incorporate the high density of information, and the trend starts to reverse, as we discuss in detail in § 3.3.3.
Scaling Inference Compute by Simply Repeating Tokens:
Many recent works around scaling test-time compute by introducing special tokens (Goyal et al., 2024a) or multiple parallel generations (Zelikman et al., 2024) have shown promising gains in reasoning tasks for language models. We test this notion with VLMs by repeating the visual input tokens (compressed to 4) multiple times to allow for more processing of key visual aspects. However, we do not observe any performance gains. This is most likely due to the fact that the downstream tasks for VLMs are not as reasoning-intensive, which demonstrates the importance of developing better token compression algorithms and potentially introducing more challenging benchmarks.
3.3.3 Scaling Laws for OCR Tasks
@@
Figure 4: Performances of various LLM size and visual token count combinations at similar inference compute. For visual reasoning tasks, at a given fixed inference cost, increasing the LLM size by decreasing the number of visual tokens improves VLM performance. However, for text recognition tasks, decreasing the number of visual tokens is detrimental to performance (§ 3.3.3).
@@
Until now, we have focused on scaling behavior for visual reasoning and understanding tasks, highlighting the key finding that using a single visual token with the maximum possible LLM parameters is the inference-optimal configuration. However, is the same valid for all tasks? VLMs have recently been applied to document reading and OCR-style tasks where a single visual token may be insufficient due to the high density of information. Unlike visual reasoning tasks, these tasks lack visual structure in the image and intuitively need more tokens to record the (generally textual) details in the image. We verify the same by fitting our scaling laws (Eq. 2) on DocVQA (Mathew et al., 2021) and TextVQA (Singh et al., 2019) benchmarks, where the tasks require mainly OCR capabilities.
@@
Figure 3(b) presents the fitted scaling law for OCR tasks. Notably, there are no significant gains in average downstream performance from increasing LLM parameters; instead, the number of visual tokens predominantly dictates the performance. This observation is reflected in the scaling law parameters, where the LLM-quality parameter $\alpha=0.029$ is nearly twice as smaller than the token quality parameter $\beta=0.048$. This trend is in stark contrast to the scaling parameters observed for visual reasoning tasks where the LLM-quality parameter ($\alpha$) was more than five times larger than the token parameter (Fig. 1(a)). This notion of visual tokens playing the significant role in OCR tasks is further echoed in Figure 4, which shows token compression weakens VLM performance despite increasing the size and capabilities of the LLM component to compensate.
4 Query-Based Token Compression
@@
Figure 5: Our query-based convolutional cross-attention (QueCC, pronounced “quick”) compression technique. User input text tokens are first processed through the LLM backbone to generate text embeddings that are then combined with the visual tokens. Within QueCC, the query-embedded visual tokens are downsampled via convolution. Next, local cross-attention is applied between the downsampled tokens and their respective visual tokens regions. The compressed tokens pass through an MLP before passing into the LLM, alongside input text tokens, for generation (§ 4).
@@
While prior work has primarily focused on moderately compressing the tokens (e.g., reducing 576 tokens to 144) while trying to match the performance of the base model (no token compression), our findings (§ 3.3.1) suggest the need for a paradigm shift. Rather than aiming for moderate token compression, new approaches should be tailored for extreme token reduction — down to 1, 4, or 16 tokens — with minimal possible degradation, as our scaling laws demonstrate that compute-optimal behavior is within this range.
Our work takes initial steps in this direction by introducing a query-based token compression strategy designed for such high-compression regimes. In cases where tokens are reduced to as few as 1, token compression based on the user’s input query becomes critical for retaining relevant information and minimizing performance reductions. In the following section, we build on existing algorithms (Li et al., 2024c), to incorporate query-based token compression. Figure 5 summarizes our query-based convolutional cross-attention (QueCC, pronounced “quick”) compression technique.
User Query Information Injection: To make our projector query-dependent, we add the text embedding of the user’s most recent query to the image embeddings from the vision encoder. We do this by taking the last hidden states prior to the LM head of the user input from the language model as the representation of the user’s overall query. The hidden state is converted into the text embedding via a linear projection and added to the image visual token embeddings. These fused tokens are later used as the query component for cross-attention. The text embedding can easily be cached for applications where the query is static or is part of a predetermined set. Even if the query varies, the text-embedding can be precalculated prior to processing the image and KV values cached and reused when processing the visual and text tokens together during generation.
Token Downsampling with Cross-Attention and Learnable Convolutions: To compress the number of visual tokens passed into the LLM, we utilize a region-based, cross-attention mechanism that downsamples the vision encoder tokens, $\mathbf{X}$, into a more information-dense form. The mechanism hinges on viewing $\mathbf{X}$ as a $\sqrt{n}\times\sqrt{n}$ grid due to the patchification of the image by the vision encoder. Li et al. (2024c; d) passes the “2D” version of $\mathbf{X}$ through a downsampling function that compresses the input by a $s^{2}$ factor where each resulting token corresponds to a $s\times s$ region in the original input. After this, cross-attention is applied independently between each downsampled token and the corresponding tokens in its $s\times s$ region. We improve on the bilinear interpolation-based downsampling techniques (Li et al., 2024c; Wang et al., 2024b) by using a learnable, depth-wise 2D convolution filter of kernel size and stride $s$, providing better expressivity.
@@ 4.1 Experimental Setup
4.2 Query-Based Convolutional Cross-Attention (QueCC) Results
@@
Table 1: Comparison of various token compression methods for VLMs at different compression rates. All models use the Vicuna-1.5 7B model as the language backbone. A denotes benchmark results for other techniques we evaluated, while best scores are bolded, and second best underlined. Our method outperforms alternatives on almost all benchmarks at extremely high compression regions (visual tokens reduced to 1 or 4) and has strong performance at lower compression rates.
Table 1 presents the results of our QueCC algorithm in comparison to previous methods, including TokenPacker (Li et al., 2024c), LLaVa-PruMerge (Shang et al., 2024), Matryoshka Multimodal Models (Cai et al., 2024), and Matryoshka Query Transformer (Hu et al., 2024), in low token regimes. We find that our method performs better than alternatives at the highest compression levels in multiple different data sets, leading to a 12% and 19% improvement in the gap between the original LLaVA-1.5 model and the next-best method on MME and MMB for the one-visual-token level. The trend continues at the four-token level, where the gap between QueCC and the next-best algorithm was reduced by 26% and 21% on MME and MMB. Our model exhibits strong performance on GQA, MME, SQA, and VQAv2 across compression rates, signaling the prospects of using the user’s query to identify and compress key visual tokens.
@@ 5.1 Token Reduction in Vision-Language Models (VLMs)
VLMs are composed of three key components: (a) a visual encoder that encodes the input images, (b) a large language model (LLM) that processes the visual tokens from the encoder along with the user text query, and (c) a projector that maps the visual tokens to the input embedding space of the LLM. Section A.1 contains additional details exploring various projector designs. Often, the number of visual tokens (576 tokens per image for CLIP-ViT-L, for instance) significantly exceeds the number of text tokens, leading to high inference costs. This disproportionate scaling of visual tokens also hinders multi-frame integration due to the limited context length of the model. In addition, inference cost is a critical factor in many real-world applications of computer vision systems. Thus, reducing the number of visual tokens processed by the language model has become an active area of research.
@@
Another approach to reducing inference cost is adaptive token processing, where the compute dedicated to certain tokens at inference is varied Jain et al. (2024). Most of these methods prune visual tokens within the LLM due to their lower attention scores compared to the prompt, system, etc., tokens (Chen et al., 2024; Wan et al., 2024), a heuristic commonly found in regular text-only LLM KV cache reduction techniques (Zhang et al., 2023; Oren et al., 2024). Finally, while we focus our paper on image-based VLMs, a host of works (Xu et al., 2024; Shen et al., 2024) discuss token compression for video processing using VLMs. We defer a discussion of these to Section A.1.
In contrast to the works discussed above that focus on developing and improving token reduction techniques, our work aims to characterize the optimal trade-off between LLM parameters and the number of visual tokens to minimize inference costs. This analysis is crucial for guiding decisions on the extent to which tokens should be reduced relative to LLM parameters given a specific target for inference cost reduction. For our inference optimal scaling laws in this work, we use TokenPacker (Li et al., 2024c) for token reduction because of its better downstream performance compared to other options (see Table 1). However, our observations from scaling laws naturally extend to any other competitive token reduction technique.
@@ 5.2 Scaling Laws and Scaling Inference Compute
Understanding how the performance of modern deep networks improves as key design factors, such as the number of parameters or training tokens, are scaled has become a focal point of research, particularly as these models continue to grow in size and complexity. Scaling laws offer crucial guidance for optimizing the architecture of such models. Notably, Kaplan et al. (2020); Hernandez et al. (2021); Hoffmann et al. (2022) do a thorough investigation into training compute-optimal language models, highlighting the need to scale pretraining tokens and parameters at the same rate. Cherti et al. (2023); Gadre et al. (2023) perform a similar study on scaling laws for CLIP (Radford et al., 2021), corroborating that performance improvements arise from increasing both parameter counts and pretraining image-caption pairs.
@@ 6 Discussion and Conclusion
In our work, we demonstrate that the optimal trade-off for VLMs inference is to use very few visual input tokens along with the largest possible LLM that fits within the budget. This result has quite important consequences. Existing works aim towards moderate reduction in token count (e.g., from 576 to 144), while trying to match the performance of the base model (no token reduction). However, our results show that the community needs to focus towards extreme token reduction (e.g., down to 1, 4 or 16 tokens), as the inference optimal regime requires very few visual input tokens. Note that although extreme token reduction can lead to a drop in performance compared to the base model, it is still better than using more tokens with a smaller LLM. The performance with very few visual tokens is poised to only improve further as we develop token reduction algorithms tailored for extreme reduction. Our work takes an initial step in this direction by proposing input query-based token reduction, as it is better to prioritize visual tokens with information relevant to the text input query, under such an extreme token compression. While our findings are focused on visual token compression at the projector level prior to passing into the LLM, we leave the compute-optimal scaling properties of adaptive token processing algorithms that operate within the LLM component for subsequent work. We hope that these critical insights from our paper will guide future research towards developing better token reduction techniques and thus inference optimal VLMs.

### v2

@@ 3.3.2 Variation in Optimal Tokens with Text Query Length
(a) Performance trend changes based on LLM sizes when varying number of input text token $Q$.
(b) Scaling law for VLM token compression and LLM model size on OCR-like tasks.
@@
Figure 3: Performance trends when shifting input text token count and benchmark family. Left: For visual reasoning tasks, as the number of text tokens increases, the impact of increasing the number of visual tokens $V$, i.e., reducing compression, becomes more apparent. Intuitively, at a large enough amount of text tokens, initial increases in visual tokens are only a minor fraction of the overall compute. Right: When the family of tasks shifts from visual reasoning to OCR/text-understanding, the trends shift: visual token count should be the prioritized instead of LLM size.
In the previous section, we observed that when the text input can be cached ($Q=0$), compute optimal inference requires the use of a single visual token paired with the largest possible LLM that fits under the inference budget. This scenario covers many practical applications, such as monitoring systems or scenarios where text input remains static. However, in interactive systems where the text input can be dynamic and long, i.e., high $Q$, the situation changes.
In Figure 3(a), we plot the average downstream error against FLOPs across different lengths of text input tokens ($Q$), with the color of the lines representing the variations in $Q$. When comparing the performance of the 7B model (solid curves) with the 4B model (dashed curves) at a high $Q$ (indicated by the green curves for each model), we observe that there is a sharp increase in error as inference FLOPs are reduced for the 7B model, particularly when visual tokens are reduced significantly. At a certain point (marked by the red dot in Fig. 3(a)), it becomes more advantageous to use the 4B model with a higher number of visual tokens rather than the 7B model with fewer tokens.
This phenomenon can be understood intuitively: as the LLM processes longer text sequences, the computational cost incurred by text tokens is already considerable. Consequently, increasing the number of visual tokens has a comparatively smaller impact on the overall inference FLOPs. Therefore, for higher text token lengths ($Q$), increasing the number of visual tokens leads to better performance without significantly increasing the computational burden. Thus, the optimal number of visual input tokens rises with an increase in $Q$. This case demonstrates the need for careful balancing of visual token count and LLM size, especially in scenarios where text inputs are long, to achieve compute-optimal performance without sacrificing accuracy.
3.4 Scaling Laws for OCR Tasks
@@
Figure 4: Performances of various LLM size and visual token count combinations with similar inference compute on two families of tasks. For many visual reasoning tasks, increasing the LLM size by decreasing the number of visual tokens improves performance. However, for text recognition tasks, decreasing the number of visual tokens is detrimental to performance.
@@
Until now, we have focused on scaling behavior for visual reasoning tasks, highlighting the key finding that using a single visual token with the maximum possible LLM parameters is the inference-optimal configuration. However, is the same valid for all tasks? VLMs have recently been applied to document reading and OCR-style tasks where a single visual token may be insufficient due to the high density of information. Unlike visual reasoning tasks, these tasks lack visual structure in the image and intuitively need more tokens to record the (generally textual) details in the image. We verify the same by fitting our scaling laws (Eq. 2) on DocVQA (Mathew et al., 2021) and TextVQA (Singh et al., 2019) benchmarks, where the tasks require mainly OCR capabilities.
@@
Figure 3(b) presents the fitted scaling law for OCR tasks. Notably, there are no significant gains in average downstream performance from increasing LLM parameters; instead, the number of visual tokens predominantly dictates the performance. This observation is reflected in the scaling law parameters, where the LLM-quality parameter $\alpha=0.029$ is nearly twice as smaller than the token quality parameter $\beta=0.048$. This trend is in stark contrast to the scaling parameters observed for visual reasoning tasks where the LLM-quality parameter ($\alpha$) was more than five times larger than the token parameter (§3.3). This notion of visual tokens playing the significant role in text-in-image recognition and understanding is further echoed in Figure 4, which shows token compression weakens VLM performance despite increasing the capabilities of the LLM component to compensate.
3.5 Generalizing Scaling Laws to other Token Compression Algorithms
We find that the takeways for our proposed scaling laws generalize across visual token compression algorithms. We fit scaling laws with VLMs utilizing LLaVA-PruMerge (Shang et al., 2024), one of the first visual token compression projectors, on $N\in\{0.5,1.8,4\}B$, and $T\in\{36,64,144,192,228,576\}$ following Section 3.2. Unlike many current projectors (Cai et al., 2024; Hu et al., 2024; Li et al., 2024c), LLaVA-Prumerge suffers massive performance drops in extreme token compression regimes, resulting from its training-free methodology. Therefore, we do not consider these conditions in its scaling laws.
When using the same $A,B,D$ values fit in Section 3.3, we find comparable $\alpha=0.069,\beta=0.008$ compared to before ($\alpha=0.077,\beta=0.015$). Similar values for $\alpha$ show that our scaling law is capable of capturing the quality of the LLM across VLM architectures and the decrease in $\beta$ shows that the PruMerge compression algorithm “weaker” than TokenPacker, which was also empirically shown in our Table 1. Fitting the scaling laws from scratch results in $\alpha=0.077,\beta=0.041$ where performance error only varies around 2×2$\times$2 × faster with LLM parameters than with tokens. Thus, even across different VLM architectures, compute-optimal inference for visual reasoning and understanding tasks continues to strongly favor the LLM parameter count, as shown in Figure 5.
Figure 5: Inference optimal scaling laws for PruMerge: When replacing the token compression algorithm, the main findings still hold: inference-optimal behavior is still to increase the LLM parameter count by reducing visual tokens in fixed compute scenarios.
4 Enhancing Extreme Token Compression with User Query
@@
Figure 6: Our query-based convolutional cross-attention compression technique. User input text tokens are first processed through the LLM backbone to generate text embeddings that are then combined with the visual tokens. Within the projector, the query-embedded visual tokens are downsampled via convolution. Next, local cross-attention is applied between the downsampled tokens and their respective visual tokens regions. The compressed tokens finally pass through an MLP into the LLM.
@@
Our scaling laws show that, for visual reasoning tasks, it is optimal to utilize the largest LLM backbone by reducing the number of visual tokens to remain within budget. Many current token compression algorithms focus on compression to 144, 64, or 36 tokens; however, the token count can be as low as 1 for many situations. Thus, improving the ability of extreme compression algorithms would enable the use of larger LLM-backbones for the same inference cost, achieving optimal inference compute usage.
Our work takes initial steps in this direction by introducing a query-based token compression strategy designed for such high-compression regimes. In cases where tokens are reduced to as few as 1, token compression based on the user’s input query becomes critical for retaining relevant information and minimizing performance reductions. In the following section, we build on existing algorithms (Li et al., 2024c), to incorporate query-based token compression. Figure 6 summarizes our query-based convolutional cross-attention (QueCC) compression technique.
User Query Information Injection: To make our projector prompt/query-dependent, we add the text embedding of the user’s most recent prompt to the image embeddings from vision encoder. We do this by taking the last hidden states prior to the LM head of the user input from the language model as the representation of the user’s overall query. The hidden state converted into the text embedding via a linear projection and added to the image visual token embeddings. These fused tokens are later used as the query component for cross-attention. The text-embedding can easily be cached for applications where the prompt is static or is part of a predetermined set. Even if the prompt varies, the text-embedding can be pre-calculated prior to processing the image and KV values cached and re-used when processing the visual and text tokens together for generation.
Token Downsampling with Cross-Attention and Learnable Convolutions: To compress the number of visual tokens passed into the LLM, we utilize a region-based, cross-attention mechanism that downsamples the vision encoder tokens, $\mathbf{X}$, into a more information-dense form. The mechanism hinges on the property that the $\mathbf{X}$ can be viewed as a $\sqrt{n}\times\sqrt{n}$ grid due to the vision encoder’s patchification of the image. Li et al. (2024c; d) passes the “2D” version of $\mathbf{X}$ through a downsampling function that compresses the input by a $s^{2}$ factor where each resulting token corresponds with a $s\times s$ region in the original input. After this, cross-attention is applied independently between each downsampled token and the corresponding tokens in its $s\times s$ region. We improve upon bilinear interpolation-based downsampling techniques (Li et al., 2024c; Wang et al., 2024b) by using a learnable depth-wise 2D convolution filter of kernel size and stride $s$, providing better expressivity.
@@ 4.1 Experimental Setup
4.2 Query-Based Convolutional Cross-Attention Results
@@
No Visual Tokens 0 $37.7$ $21.0$ $697.8$ $45.4$ $63.6$ $41.7$ $44.4$ $41.0$
Table 1: Comparison of various token compression methods for VLMs at different compression rates. All models use the Vicuna-1.5 7B model as the language backbone. A denotes benchmark results for other techniques we evaluated, while best scores are bolded, and second best underlined. Our updated method outperforms alternatives, including its predecessor, on almost all benchmarks at extremely high compression regions (visual tokens reduced to 1 or 4), highlighting the benefit of utilizing the user query in the compression algorithm for these situations.
Table 1 presents the results of our query-based algorithm in comparison to previous methods, including TokenPacker (Li et al., 2024c), LLaVA-PruMerge (Shang et al., 2024), Matryoshka Multimodal Models (Cai et al., 2024), and Matryoshka Query Transformer (Hu et al., 2024), in low token regimes. We find that our query based token compression (QueCC) performs better than alternatives at the highest levels of compression on multiple different datasets. At the one-token level, our method outperforms other methods on six of the eight datasets and is able to mitigate some of the shortcomings of the original TokenPacker by reducing its gap between vanilla LLaVA-1.5 on VizWiz by $34\%$ and ${\sim}12\%$ on both POPE and MMBench. The trend continues at the four-token level. Our model also exhibits strong performance on GQA, MME, SQA, and VQAv2 across compression rates, signaling the prospects of using the user’s query to identify key tokens.
@@ 5.1 Token Reduction in Vision-Language Models (VLMs)
VLMs are composed of three key components: (a) a visual encoder that encodes the input images, (b) a language model (LM) that processes the visual tokens from the encoder along with the user text query, and (c) a projector that maps the visual tokens to the input embedding space of the LM. Appendix A.1 contains additional details exploring various projector designs. Often, the number of visual tokens (576 tokens per image for CLIP-ViT-L, for instance) significantly exceeds the number of text tokens, leading to high inference costs. This disproportionate scaling of visual tokens also hinders multi-frame integration due to the limited context length of the model. Inference cost is a critical factor in many real world applications of computer vision systems. Thus, reducing the number of visual tokens processed by the language model has become an active area of research.
@@
Another approach to reducing inference cost is adaptive token processing, where the compute dedicated to certain tokens at inference is varied Jain et al. (2024). Many methods prune visual tokens within the LLM due to their lower attention scores compared to the prompt, system, etc., tokens (Chen et al., 2024; Wan et al., 2024), a heuristic commonly found in regular text-only LLM KV cache reduction techniques (Zhang et al., 2023; Oren et al., 2024). Finally, while we focus our paper on image-based VLMs, a host of works (Xu et al., 2024; Shen et al., 2024) discuss token compression for video processing using VLMs. We defer a discussion of these to Appendix A.1.
@@ 5.2 Scaling Laws and Scaling Inference Compute
Understanding how the performance of modern deep networks shifts as key design factors, such as the number of parameters or training tokens, are scaled has become a focal point of research, particularly as models continue to grow in size and complexity. Scaling laws offer crucial guidance for optimizing the architecture of such models. Notably, Kaplan et al. (2020); Hernandez et al. (2021); Hoffmann et al. (2022) do a thorough investigation into training compute-optimal language models, highlighting the need to scale pretraining tokens and parameters at the same rate. Cherti et al. (2023); Gadre et al. (2023) perform a similar study on scaling laws for CLIP (Radford et al., 2021), corroborating that performance improvements arise from increasing both parameter counts and pretraining image-caption pairs.
@@ 6 Discussion and Conclusion
In our work, we demonstrate that the optimal trade-off for VLMs inference is to use very few visual input tokens along with the largest possible LLM that fits within the budget. This result has quite important consequences. Existing works aim towards moderate reduction in token count (e.g., from 576 to 144), while trying to match the performance of the base model (no token reduction). However, our results show that the community needs to focus towards extreme token reduction (e.g., down to 1, 4 or 16 tokens), as the inference optimal regime requires very few visual input tokens. Note that although extreme token reduction can lead to a drop in performance compared to the base model, it is still better than using more tokens with a smaller LLM. The performance with very few visual tokens is poised to only improve further as we develop token reduction algorithms tailored for extreme reduction. While our findings are focused on visual token compression at the projector level prior to passing into the LLM, we leave the compute-optimal scaling properties of adaptive token processing algorithms that operate within the LLM component for subsequent work. We hope that these critical insights from our paper will guide future research towards developing better token reduction techniques and thus inference optimal VLMs.